# Moser’s twist theorem revisited

### v2

@@ 4.3. Growth of the distortions
To prove Item (2), we write
$\Theta(v_{3},v_{4})K^{2}(j|v_{1},\ldots,v_{4})=$ $\sum_{i=0}^{j-1}\left(K^{1}(F^{i}(v_{3},v_{4}))-\frac{\Theta(F^{i}(v_{3},v_{4}))}{\Theta(F^{i}(v_{1},v_{2}))}K^{1}(F^{i}(v_{1},v_{2}))\right)$
$-\sum_{i=0}^{j-1}\left(\frac{\Theta(v_{3},v_{4})}{\Theta(v_{1},v_{2})}-\frac{\Theta(F^{i}(v_{3},v_{4}))}{\Theta(F^{i}(v_{1},v_{2}))}\right)K^{1}(F^{i}(v_{1},v_{2})).$
@@
$\frac{\Theta(F^{i}(v_{1},v_{2}))}{\Theta(F^{i}(v_{3},v_{4}))}\left(\frac{\Theta(v_{3},v_{4})}{\Theta(v_{1},v_{2})}-\frac{\Theta(F^{i}(v_{3},v_{4}))}{\Theta(F^{i}(v_{1},v_{2}))}\right)$
$\leq$ $\left|\frac{\Theta(v_{3},v_{4})}{\Theta(F^{i}(v_{3},v_{4}))}-\frac{\Theta(v_{1},v_{2})}{\Theta(F^{i}(v_{1},v_{2}))}\right|\frac{\Theta(F^{i}(v_{1},v_{2}))}{\Theta(v_{1},v_{2})}.$
By the definition of $\Theta(v_{i},v_{j})$, we have
$\frac{\Theta(F^{i}(v_{1},v_{2}))}{\Theta(v_{1},v_{2})}\leq e^{\frac{\varepsilon}{40}R},\quad\left|\frac{\Theta(v_{3},v_{4})}{\Theta(F^{i}(v_{3},v_{4}))}-\frac{\Theta(v_{1},v_{2})}{\Theta(F^{i}(v_{1},v_{2}))}\right|\lesssim e^{\frac{\varepsilon}{40}R}\tilde{\mathbb{K}}_{r}^{1}(R,R).$
@@
$\frac{\Theta(F^{i}(v_{1},v_{2}))}{\Theta(F^{i}(v_{3},v_{4}))}\left(\frac{\Theta(v_{3},v_{4})}{\Theta(v_{1},v_{2})}-\frac{\Theta(F^{i}(v_{3},v_{4}))}{\Theta(F^{i}(v_{1},v_{2}))}\right)\lesssim\varepsilon e^{\frac{\varepsilon}{10}R}q_{r}q_{R}^{-1}\leq\varepsilon(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
[…]
A direct calculation shows (see [11, Proposition 2.3])
Lemma 4.5.
Given $(v_{1},v_{2},v_{3},v_{4})\in(\kappa,r,s)_{II}$, we have
$K^{2}(v_{1},\ldots,v_{4})\lesssim\nabla^{2}(v_{1},\ldots,v_{4})+\Theta(v_{1},\ldots,v_{4}).$
@@
$\kappa^{1}(r;R,\kappa):={\sup}\left\{\frac{|K^{1}(q_{r}|v_{1},v_{2})|}{\sum_{i=0}^{q_{r}-1}\Theta(F^{i}(v_{1},v_{2}))}\ \big{|}\ (v_{1},v_{2})\in(\kappa,R)_{II}\right\}.$
@@
Proof   It suffices to prove $q_{M}\leq q_{m+[\frac{m}{2}]}$. Due to the requirement on $\delta$, a direct calculation shows for $m\geq 20$,
@@
[…]
Based on Lemma 4.4(1), under the assumption (19), there holds $\tilde{\mathbb{K}}_{r}^{1}(\kappa,R)\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}$. It follows that for $q_{R}^{1-\frac{\varepsilon}{2}}\sim q_{r}$,
$\tilde{\mathbb{K}}_{r}^{1}(\kappa,R)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}},$
where $r\ll R$ due to the requirement $q_{R}^{1-\frac{\varepsilon}{2}}\sim q_{r}$. In the following, we will show that with the help of the assumption (20), $\tilde{\mathbb{K}}_{r}^{1}(\kappa,R)$ is still exponentially small even for $r=R$, i.e.
$\tilde{\mathbb{K}}_{R}^{1}(\kappa,R)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{3}}.$
@@
(30) $\kappa^{1}(r;R,\kappa)\lesssim(\frac{9}{10})^{\frac{\varepsilon r}{2}}.$
In particular,
$\kappa^{1}(R;R,\kappa)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}},\quad\tilde{\mathbb{K}}_{R}^{1}(\kappa,R)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{3}}.$
Proof   We prove this result by contradiction. We assume (30) does not hold. For fixed $R$ and $\kappa$, we write $\kappa^{1}(r)$ instead of $\kappa^{1}(r;R,\kappa)$ for simplicity.
@@
(31) $\tilde{\mathbb{K}}_{R}^{1}(\kappa,R)\leq 2\kappa^{1}(r).$
@@
(32) ${K^{1}(j|v_{1},v_{2})\leq\kappa^{1}(r)\sum_{i=0}^{j-1}\Theta(F^{i}(v_{1},v_{2}))+\tilde{\mathbb{K}}_{r}^{1}(\kappa,R).}$
If (30) does not hold, it follows from Lemma 4.4(1) that
$K^{1}(j|v_{1},v_{2})\leq 2\kappa^{1}(r),$
@@
(33) $\kappa^{1}(R)\lesssim\kappa^{1}(r)^{2}+(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
@@
$K^{1}(q_{R}|v^{\prime}_{1},v^{\prime}_{2})\leq 0.$
Then for each $w^{\prime}>0$, we have
@@
$i^{\prime}:=\left\{\begin{array}[]{ll}i-i_{0},\quad i_{0}\leq i<q_{R},\\ i-i_{0}+q_{R},\quad 0\leq i<i_{0}.\end{array}\right.$
Then $F^{i_{0}+i}(v_{1},v_{2})$ and $F^{i}(v^{\prime}_{1},v^{\prime}_{2})$ are disjoint. Different from the notation of (17), we denote
$E(j|v_{1},v_{2}):=\frac{\Theta(F^{j}(v_{1},v_{2}))}{\Theta(v_{1},v_{2})},$
$\Upsilon:=\frac{\Theta(F^{i^{\prime}}(v^{\prime}_{1},v^{\prime}_{2}))}{\Theta(v^{\prime}_{1},v^{\prime}_{2})}\cdot\frac{\Theta(F^{i_{0}}(v_{1},v_{2}))}{\Theta(F^{i}(v_{1},v_{2}))}$
@@
(35) $\Upsilon=\left\{\begin{array}[]{ll}\frac{E(i^{\prime}|v^{\prime}_{1},v^{\prime}_{2})}{E(i^{\prime}|F^{i_{0}}(v_{1},v_{2}))},\quad i_{0}\leq i<q_{R},\\ \frac{E(i^{\prime}|v^{\prime}_{1},v^{\prime}_{2})}{E(i^{\prime}|F^{i_{0}-q_{R}}(v_{1},v_{2}))\cdot E(q_{R}|F^{-q_{R}}(v_{1},v_{2}))},\quad 0\leq i<i_{0}.\end{array}\right.$
@@
$\Upsilon\in\left(\exp(-3\tilde{\mathbb{K}}_{R}^{1}(\kappa,R)),\exp(3\tilde{\mathbb{K}}_{R}^{1}(\kappa,R))\right).$
@@
$\Upsilon\in\left(\exp(-6\kappa^{1}(r)),\exp(6\kappa^{1}(r))\right).$
@@
$w^{\prime}=\exp(-6\kappa^{1}(r))\frac{\Theta(F^{i_{0}}(v_{1},v_{2}))}{\Theta(v^{\prime}_{1},v^{\prime}_{2})},$
@@
(36) $\exp(-12\kappa^{1}(r))\leq w^{\prime}\frac{\Theta(F^{i^{\prime}}(v^{\prime}_{1},v^{\prime}_{2}))}{\Theta(F^{i}(v_{1},v_{2}))}\leq 1.$
@@
$\Xi_{1}:=\sum_{p}(1-w_{p})K^{1}(q_{r}|F^{b_{p}}(v_{1},v_{2})),\quad\Xi_{2}:=w^{\prime}\sum_{p}B_{p},$
(38) $B_{p}=\left\{\begin{array}[]{ll}\Theta(F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2}))K^{2}(q_{r}|F^{b_{p}}(v_{1},v_{2}),F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2})),\quad j\geq q_{r},\\ \Theta(F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2}))K^{2}(j|F^{b_{p}}(v_{1},v_{2}),F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2})),\quad j<q_{r}.\end{array}\right.$
$w_{p}:=w^{\prime}\frac{\Theta(F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2}))}{\Theta(F^{b_{p}}(v_{1},v_{2}))},$
@@
$\Xi_{1}\leq(1-\exp(-12\kappa^{1}(r)))\kappa^{1}(r)\sum_{p}\Theta(F^{p}(v_{1},v_{2})).$
By Lemma 4.4(2), we write $\Xi_{2}\leq\Xi_{21}+\Xi_{22}$ where
$\Xi_{21}:=w^{\prime}\tilde{\mathbb{K}}_{r}^{1}(\kappa,R)\sum_{i=0}^{q_{R}-1}\Theta(F^{i}(v^{\prime}_{1},v^{\prime}_{2})),$
$\Xi_{22}:=w^{\prime}\sum_{i=0}^{q_{R}-1}\Theta(F^{i^{\prime}}(v^{\prime}_{1},v^{\prime}_{2}))K^{2}(F^{i}(v_{1},v_{2}),F^{i^{\prime}}(v^{\prime}_{1},v^{\prime}_{2})).$
@@
$\Xi_{21}\lesssim w^{\prime}\varepsilon(\frac{9}{10})^{\frac{\varepsilon R}{2}}\sum_{i=0}^{q_{R}-1}\Theta(F^{i}(v^{\prime}_{1},v^{\prime}_{2}))\leq\varepsilon(\frac{9}{10})^{\frac{\varepsilon R}{2}}\sum_{i=0}^{q_{R}-1}\Theta(F^{i}(v_{1},v_{2})).$
Note that $e^{\frac{\varepsilon}{10}R}>1$. By Lemma 4.5, Lemma 5.1 and (36),
$\Xi_{22}\lesssim e^{\frac{\varepsilon}{10}R}q_{R}^{-\frac{\varepsilon}{2}}w^{\prime}\sum_{i=0}^{q_{R}-1}\Theta(F^{i}(v^{\prime}_{1},v^{\prime}_{2}))\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}}\sum_{i=0}^{q_{R}-1}\Theta(F^{i}(v_{1},v_{2})).$
@@
$\kappa^{1}(R,R;\kappa)-(1-\exp(-12\kappa^{1}(r)))\kappa^{1}(r)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}},$
which from the assumption implies $\kappa^{1}(R)<\kappa^{1}(r)$ and $\kappa^{1}(r)\to 0$ as $r\to\infty$ (up to a subsequence). Moreover, we have (33). This completes the proof of Claim 2.
@@
$\frac{\kappa^{1}(r)^{2}}{(\frac{9}{10})^{\varepsilon r}}\to+\infty.$
@@
$\frac{\kappa^{1}(r)^{2}}{(\frac{9}{10})^{\frac{\varepsilon r}{2}(\frac{3}{4})^{n}}}\to+\infty,$
which contradicts the fact that $\kappa^{1}(r)\to 0$ as $r\to\infty$. Using $r<R<\frac{3}{2}r$ again, we have
$\kappa^{1}(R)\lesssim(\frac{9}{10})^{\frac{\varepsilon r}{2}}+(\frac{9}{10})^{\frac{\varepsilon R}{2}}\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{3}}.$
@@
$\tilde{\mathbb{K}}_{R}^{1}(\kappa,R)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{3}}.$
[…]
@@
Based on Lemma 4.3, we obtain $\tilde{\mathbb{K}}_{\bar{N}(\kappa)}^{0}(\kappa)\lesssim 1$. It follows that for $0\leq R\leq\tilde{N}(\kappa)$, the factor $e^{\frac{\varepsilon R}{40}}$ can be replaced by a constant independent of $\kappa$. By repeating the argument in Lemma 4.4, we obtain
$\tilde{\mathbb{K}}_{r}^{1}(\kappa,R)\lesssim\varepsilon q_{r}q_{R}^{-1}.$
@@
$\tilde{\mathbb{K}}_{r}^{1}(\kappa,r)\lesssim q_{r}^{-\frac{\varepsilon}{3}},$

### v3

@@ 4.4.2. Exponential decay of $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)$ with $r\ll R$
To prove Item (2), we write from (36) that
$\Theta_{2}(v_{3},v_{4})K^{2}(j|v_{1},\ldots,v_{4})=$ $\sum_{i=0}^{j-1}\left(K^{1}(F^{i}(v_{3},v_{4}))-\frac{\Theta_{2}(F^{i}(v_{3},v_{4}))}{\Theta_{2}(F^{i}(v_{1},v_{2}))}K^{1}(F^{i}(v_{1},v_{2}))\right)$
$-\sum_{i=0}^{j-1}\left(\frac{\Theta_{2}(v_{3},v_{4})}{\Theta_{2}(v_{1},v_{2})}-\frac{\Theta_{2}(F^{i}(v_{3},v_{4}))}{\Theta_{2}(F^{i}(v_{1},v_{2}))}\right)K^{1}(F^{i}(v_{1},v_{2})).$
@@
$\frac{\Theta_{2}(F^{i}(v_{1},v_{2}))}{\Theta_{2}(F^{i}(v_{3},v_{4}))}\left(\frac{\Theta_{2}(v_{3},v_{4})}{\Theta_{2}(v_{1},v_{2})}-\frac{\Theta_{2}(F^{i}(v_{3},v_{4}))}{\Theta_{2}(F^{i}(v_{1},v_{2}))}\right)$
$\leq$ $\left|\frac{\Theta_{2}(v_{3},v_{4})}{\Theta_{2}(F^{i}(v_{3},v_{4}))}-\frac{\Theta_{2}(v_{1},v_{2})}{\Theta_{2}(F^{i}(v_{1},v_{2}))}\right|\frac{\Theta_{2}(F^{i}(v_{1},v_{2}))}{\Theta_{2}(v_{1},v_{2})}.$
By the definition of $\Theta_{2}(v_{i},v_{j})$, we have
$\frac{\Theta_{2}(F^{i}(v_{1},v_{2}))}{\Theta_{2}(v_{1},v_{2})}\leq e^{\frac{\varepsilon}{40}R},\quad\left|\frac{\Theta_{2}(v_{3},v_{4})}{\Theta_{2}(F^{i}(v_{3},v_{4}))}-\frac{\Theta_{2}(v_{1},v_{2})}{\Theta_{2}(F^{i}(v_{1},v_{2}))}\right|\lesssim e^{\frac{\varepsilon}{40}R}\widetilde{\mathbf{K}}_{r}^{1}(\phi(R),R).$
@@
$\frac{\Theta_{2}(F^{i}(v_{1},v_{2}))}{\Theta_{2}(F^{i}(v_{3},v_{4}))}\left(\frac{\Theta_{2}(v_{3},v_{4})}{\Theta_{2}(v_{1},v_{2})}-\frac{\Theta_{2}(F^{i}(v_{3},v_{4}))}{\Theta_{2}(F^{i}(v_{1},v_{2}))}\right)\lesssim\varepsilon e^{\frac{\varepsilon}{10}R}q_{r}q_{R}^{-1}\leq\varepsilon(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
By Lemma 4.2 and the assumption on $\nabla^{1}(\kappa,r)$, we know that $\frac{K^{1}(F^{i}(v_{1},v_{2}))}{\Theta_{2}(F^{i}(v_{1},v_{2}))}$ is bounded. Combining with (36), we obtain Item (2). ∎
4.4.3. Exponential decay of $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)$ with $r=R$
@@
$\Delta(r;R,\kappa):={\sup}\left\{\frac{|K^{1}(q_{r}|v_{1},v_{2})|}{\sum_{i=0}^{q_{r}-1}\Theta_{2}(F^{i}(v_{1},v_{2}))}\ \big|\ (v_{1},v_{2})\in(\kappa,R)_{II}\right\}.$
To improve the estimate (42), the key point is to refine the estimate for (39) to an exponential one. Essentially, this improvement is guaranteed by condition (28). Observe that in (39) one has (see (38))
(43) $\sum_{i=0}^{q_{r}-1}\Theta_{2}\bigl(F^{i}(v_{1},v_{2})\bigr)\leq 2.$
Thus it suffices to prove that $\Delta(r;R,\kappa)$ decays exponentially.
@@
Proof.
It suffices to prove $q_{M}\leq q_{m+[\frac{m}{2}]}$. Due to the requirement on $\delta$, a direct calculation shows for $m\geq 20$,
@@
which implies $q_{M}\leq q_{m+[\frac{m}{2}]}$. ∎
Based on Lemma 4.6(1), under the assumption (27), there holds $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)\lesssim\varepsilon e^{\frac{\varepsilon}{40}R}q_{r}q_{R}^{-1}$. It follows that for $q_{R}^{1-\frac{\varepsilon}{2}}\sim q_{r}$,
(44) $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}},$
where $r\ll R$ due to the requirement $q_{R}^{1-\frac{\varepsilon}{2}}\sim q_{r}$. In the following, we will show that with the help of the assumption (28), $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)$ is still exponentially small even for $r=R$, i.e.
$\widetilde{\mathbf{K}}_{R}^{1}(\kappa,R)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
@@
(45) $\Delta(r;R,\kappa)\lesssim(\frac{9}{10})^{\frac{\varepsilon r}{2}}.$
Moreover,
$\Delta(R;R,\kappa)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}},\quad\widetilde{\mathbf{K}}_{R}^{1}(\kappa,R)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
Proof.
We prove this result by contradiction. We assume (45) does not hold. For fixed $R$ and $\kappa$, we write $\Delta(r)$ instead of $\Delta(r;R,\kappa)$ for simplicity.
@@
(46) $\widetilde{\mathbf{K}}_{R}^{1}(\kappa,R)\leq 3\Delta(r).$
@@
(47) $K^{1}(j|v_{1},v_{2})\leq\Delta(r)\sum_{i=0}^{j-1}\Theta_{2}(F^{i}(v_{1},v_{2}))+\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R).$}
If (45) does not hold, it follows from (44) that the growth of $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)$ is negligible compared with $\Delta(r)$. Together with (43) this yields
$K^{1}(j|v_{1},v_{2})\leq 3\Delta(r),$
@@
(48) $\Delta(R)\lesssim\Delta(r)^{2}+(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
@@
$K^{1}(q_{R}|v^{\prime}_{1},v^{\prime}_{2})\leq 0,$
otherwise, we exchange the roles of $v^{\prime}_{1}$ and $v^{\prime}_{2}$. Then for each $w^{\prime}>0$, we have
@@
$i^{\prime}:=\left\{\begin{array}[]{ll}\hskip-3.99994pti-i_{0},\quad i_{0}\leq i<q_{R},\\ \hskip-3.99994pti-i_{0}+q_{R},\quad 0\leq i<i_{0}.\end{array}\right$.
Then $F^{i_{0}+i}(v_{1},v_{2})$ and $F^{i}(v^{\prime}_{1},v^{\prime}_{2})$ are disjoint. We denote
$Y(j|v_{1},v_{2}):=\frac{\Theta_{2}(F^{j}(v_{1},v_{2}))}{\Theta_{2}(v_{1},v_{2})},$
$\Upsilon:=\frac{\Theta_{2}(F^{i^{\prime}}(v^{\prime}_{1},v^{\prime}_{2}))}{\Theta_{2}(v^{\prime}_{1},v^{\prime}_{2})}\cdot\frac{\Theta_{2}(F^{i_{0}}(v_{1},v_{2}))}{\Theta_{2}(F^{i}(v_{1},v_{2}))}$
@@
(50) $\Upsilon=\left\{\begin{array}[]{ll}\hskip-3.99994pt\frac{Y(i^{\prime}|v^{\prime}_{1},v^{\prime}_{2})}{Y(i^{\prime}|F^{i_{0}}(v_{1},v_{2}))},\quad i_{0}\leq i<q_{R},\\ \hskip-3.99994pt\frac{Y(i^{\prime}|v^{\prime}_{1},v^{\prime}_{2})}{Y(i^{\prime}|F^{i_{0}-q_{R}}(v_{1},v_{2}))\cdot Y(q_{R}|F^{-q_{R}}(v_{1},v_{2}))},\quad 0\leq i<i_{0}.\end{array}\right$.
@@
$\Upsilon\in\left(\exp(-3\widetilde{\mathbf{K}}_{R}^{1}(\kappa,R)),\exp(3\widetilde{\mathbf{K}}_{R}^{1}(\kappa,R))\right).$
@@
$\Upsilon\in\left(\exp(-9\Delta(r)),\exp(9\Delta(r))\right).$
@@
$w^{\prime}=\exp(-9\Delta(r))\frac{\Theta_{2}(F^{i_{0}}(v_{1},v_{2}))}{\Theta_{2}(v^{\prime}_{1},v^{\prime}_{2})},$
@@
(51) $\exp(-18\Delta(r))\leq w^{\prime}\frac{\Theta_{2}(F^{i^{\prime}}(v^{\prime}_{1},v^{\prime}_{2}))}{\Theta_{2}(F^{i}(v_{1},v_{2}))}\leq 1.$
@@
$\Xi_{1}:=\sum_{p}\left(1-w^{\prime}\frac{\Theta_{2}(F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2}))}{\Theta_{2}(F^{b_{p}}(v_{1},v_{2}))}\right)K^{1}(q_{r}|F^{b_{p}}(v_{1},v_{2})),\quad\Xi_{2}:=w^{\prime}\sum_{p}B_{p},$
(53) $B_{p}:=\left\{\begin{array}[]{ll}\hskip-3.99994pt\Theta_{2}(F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2}))K^{2}(q_{r}|F^{b_{p}}(v_{1},v_{2}),F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2})),\quad j\geq q_{r},\\ \hskip-3.99994pt\Theta_{2}(F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2}))K^{2}(j|F^{b_{p}}(v_{1},v_{2}),F^{b^{\prime}_{p}}(v^{\prime}_{1},v^{\prime}_{2})),\quad j<q_{r}.\end{array}\right$.
@@
$\Xi_{1}\leq(1-\exp(-18\Delta(r)))\Delta(r)\sum_{p}\Theta_{2}(F^{p}(v_{1},v_{2})).$
By Lemma 4.6(2), we write
(54) $\Xi_{2}\leq\Xi_{21}+\Xi_{22},$
where
$\Xi_{21}:=w^{\prime}\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)\sum_{i=0}^{q_{R}-1}\Theta_{2}(F^{i}(v^{\prime}_{1},v^{\prime}_{2})),$
$\Xi_{22}:=w^{\prime}\sum_{i=0}^{q_{R}-1}\Theta_{2}(F^{i^{\prime}}(v^{\prime}_{1},v^{\prime}_{2}))K^{2}(F^{i}(v_{1},v_{2}),F^{i^{\prime}}(v^{\prime}_{1},v^{\prime}_{2})).$
@@
$\Xi_{21}\lesssim w^{\prime}\varepsilon(\frac{9}{10})^{\frac{\varepsilon R}{2}}\sum_{i=0}^{q_{R}-1}\Theta_{2}(F^{i}(v^{\prime}_{1},v^{\prime}_{2}))\leq\varepsilon(\frac{9}{10})^{\frac{\varepsilon R}{2}}\sum_{i=0}^{q_{R}-1}\Theta_{2}(F^{i}(v_{1},v_{2})).$
Note that $e^{\frac{\varepsilon}{10}R}>1$. By Lemma 4.5, Lemma 4.6(3), (28) and (51),
$\Xi_{22}\lesssim e^{\frac{\varepsilon}{10}R}q_{R}^{-\frac{\varepsilon}{3}}w^{\prime}\sum_{i=0}^{q_{R}-1}\Theta_{2}(F^{i^{\prime}}(v^{\prime}_{1},v^{\prime}_{2}))\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}}\sum_{i=0}^{q_{R}-1}\Theta_{2}(F^{i}(v_{1},v_{2})).$
@@
$\Delta(R)-(1-\exp(-18\Delta(r)))\Delta(r)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}},$
which from the assumption implies $\Delta(R)<\Delta(r)$ and $\Delta(r)\to 0$ as $r\to\infty$ (up to a subsequence). Moreover, we have (48). This completes the proof of Claim 2.
@@
$\frac{\Delta(r)^{2}}{(\frac{9}{10})^{\varepsilon r}}\to+\infty.$
@@
$\frac{\Delta(r)^{2}}{(\frac{9}{10})^{\frac{\varepsilon r}{2}(\frac{3}{4})^{n}}}\to+\infty,$
which contradicts the fact that $\Delta(r)\to 0$ as $r\to\infty$. It follows that
$\Delta(r)\lesssim(\frac{9}{10})^{\frac{\varepsilon r}{2}},$
which contradicts our assumption at the beginning of the proof.
Using $r<R<\frac{3}{2}r$ again, we have
$\Delta(R)\lesssim(\frac{9}{10})^{\varepsilon r}+(\frac{9}{10})^{\frac{\varepsilon R}{2}}\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
@@
$\widetilde{\mathbf{K}}_{R}^{1}(\kappa,R)\lesssim(\frac{9}{10})^{\frac{\varepsilon R}{2}}.$
This completes the proof of Lemma 4.8. ∎
@@
Based on Lemma 4.3, we obtain $\widetilde{\mathbf{K}}_{\bar{N}(\kappa)}^{0}(\kappa)\lesssim 1$. It follows that for $0\leq R\leq\tilde{N}(\kappa)$, the factor $e^{\frac{\varepsilon R}{40}}$ can be replaced by a constant independent of $\kappa$. By repeating the argument in Lemma 4.6, we obtain
$\widetilde{\mathbf{K}}_{r}^{1}(\kappa,R)\lesssim\varepsilon q_{r}q_{R}^{-1}.$
@@
$\widetilde{\mathbf{K}}_{r}^{1}(\kappa,r)\lesssim q_{r}^{-\frac{\varepsilon}{3}},$